A sample of 25 patients in a doctor's office showed that they had to wait an average of 35 minutes
with a standard deviation of 10 minutes before they could see the doctor. Provide a 98%
confidence interval estimate for the average waiting time of all the patients who visit this doctor.
Assume the population of waiting times is normally distributed.
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 35 minutes

sample standard deviation = s = 10 minutes

sample size = n = 25

Degrees of freedom = df = n - 1 = 25 - 1 = 24

At 98% confidence level

= 1 - 98%

=1 - 0.98 =0.02

/2 = 0.01

t/2,df = t0.01,24 = 2.492

Margin of error = E = t/2,df * (s /n)

= 2.492 * (10 / 25)

Margin of error = E = 4.98

The 98% confidence interval estimate of the population mean is,

  ± E  

= 35  ± 4.98

=( 30.02, 39.98 )